Your client’s municipal bond portfolio is now worth $540,000. He originally invested $430,000 seven years ago (equals his cost basis). His marginal federal ordinary income tax rate is 40% and his state rate is 6%. Your client’s state exempts muni-bond interest but not realized gains from sales. The tax rate on long-term capital gains is currently 25%. He asks you to sell the entire portfolio to fund the purchase of a second home. What are the tax consequences for doing so?

A. $0

B. $27,500

C. $34,100

D. $50,600
Answer #1

gain from selling portfolio = $540000-$430000
=110000

long term capital gain tax = 110000*0.25 = 27500
state tax = 110000*0.06 = $ 6600
total tax consequnces = $34100

Hence option C is correct.